Points B, O and Y are collinear on BY, and BO:OY =5/8.
B is located at (4,2)
O is located at (57/8, -3), and Y is located at (x,y).
Determine the values of x and y.
Set up the equation below to provide of how you determined the values of x and y

Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

Formula for internal division
P={[(mx2+nx1)/(m+n)],[(my2+ny1)/(m+n)]}
where x1 y1 & x2y2 are the two points and x y is thepoint which divides in the ratio of m:n
57/8 = (4*8 +5*x)/5+8
57/8 = (4*8 +5*x)/13
13*57/8 = 32 +5x
5x = 60.625
x=12.125 = 97/8
-3 = (2*8+5y)/13
-39 = 16+5y
-55= 5y
y= -11
.



Answer by greenestamps(13200)



Find the x and y coordinates of Y separately, using the coordinates of B and O and the given ratio.

The difference between the x coordinates of B and O is 57/8-4 = 57/8-32/8 = 25/8. Since the given ratio is 5:8, the difference between the x coordinates of O and Y is (8/5)*(25/8) = 5; so the x coordinate of Y is 57/8+5 = 57/8+40/8 = 97/8.

The difference between the y coordinates of B and O is -3-2 = -5. Since the given ratio is 5:8, the difference between the y coordinates of O and Y is (8/5)*(-5) = -8; so the y coordinate of Y is -3-8 = -11.

ANSWER: Y is (x,y) = (97/8,-11)